OpenStudy (anonymous):

Two sides of a triangle have lengths 6 and 17. Which expression describes the length of the third side?

OpenStudy (icalibear):

6^2 + 17^2 = x^2 given that the third side is "x" and the formula for this is a^2 + b^2 = c^2.

OpenStudy (anonymous):

a^2+b^2=c^2 therefore 6^2+17^2=c^2 36+289=c^2 c^2=325